OpenStudy (anonymous):

Solve the system by substitution. -x - y - z = -8 -4x + 4y + 5z + 7 2x + 2z + 4

OpenStudy (anonymous):

I've done some of my work already. I'm just stuck now

OpenStudy (anonymous):

http://www.youtube.com/watch?v=g6FQBfIwf3w

OpenStudy (anonymous):

hope this helps! :) @nicki96

OpenStudy (anonymous):

i understand how to do it, im just not sure if I did one of them right

OpenStudy (aripotta):

tell us what you got then

OpenStudy (anonymous):

ill give you the answers and you check if they are right :)

OpenStudy (anonymous):

do you want me to show you the work ive done so far?

OpenStudy (aripotta):

sure

OpenStudy (anonymous):

can you please check your equations again..they're not correct

OpenStudy (anonymous):

the second two dont have equal signs

OpenStudy (anonymous):

since i have to use substitution i solved for y in the first one and got x = -y - x + 8

OpenStudy (anonymous):

I meant to put = instead of the + sign, sorry

OpenStudy (anonymous):

okay so you should have gotten x=3, y=6, and z=-1 the rest Ari will help with :D

OpenStudy (anonymous):

I solved for x in the first one so I could substitute it in for x in the other equations

OpenStudy (anonymous):

okay thanks yummydum

OpenStudy (aripotta):

x = -y - x + 8 i think you meant x = -y - z + 8

OpenStudy (aripotta):

so now we substitute that into the other two equations

OpenStudy (anonymous):

yeah that's what I meant

OpenStudy (aripotta):

-4( -y - z + 8) + 4y + 5z = 7

OpenStudy (anonymous):

okay for the second one I substituted it in and got 8y + 9z = 39

OpenStudy (anonymous):

Then for the third equation I got stuck on it

OpenStudy (aripotta):

2(-y - z + 8) + 2z = 4

OpenStudy (anonymous):

i got -2y -2z + 16 + 2z = 4

OpenStudy (anonymous):

But doesn't it cancel out z ?

OpenStudy (aripotta):

yes it does

OpenStudy (anonymous):

so you would have to solve for y?

OpenStudy (aripotta):

you have to subtract 16 on both sides first

OpenStudy (anonymous):

y = 6 is what I got

OpenStudy (aripotta):

yup

OpenStudy (anonymous):

then would you plug it into the second equation?

OpenStudy (aripotta):

so substitute that into 8y + 9z = 39 to solve for z

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i got z = -1

OpenStudy (aripotta):

yup. so now plug those numbers into any of the original equations to solve for x

OpenStudy (anonymous):

would you then plug both y and z into one of the original three equations?

OpenStudy (aripotta):

yes lol

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

i got x = 3, so the answer would be (3,6,-1) ?

OpenStudy (aripotta):

yup

OpenStudy (anonymous):

yay! thank you for your help!

OpenStudy (aripotta):

no problem :)